OpenStudy (anonymous):

The distance d between two points (x, y) and (a, b) is given by d =( (x - a)^2 + (y - b)^2)^0.5 . Use the method of Lagrange Multipliers to find the points on the circle x^2 - 2x + y^2 - 4y =0 which are nearest to and furthest from the point (-2,-1). (Hint: Consider d2 subject to the constraint.)

OpenStudy (anonymous):

i think lagrange better solve this one

OpenStudy (anonymous):

oh he's offline

OpenStudy (anonymous):

sorry dont know this

OpenStudy (anonymous):

Anybody can help me?!

OpenStudy (dumbcow):

i solved it but i didn't use Lagrange Multipliers. solve circle equation for y \[y = 2 \pm \sqrt{5-(x-1)^{2}}\] Then use distance formula to obtain function in terms of x representing the distance from a point on circle to point (-2,-1) \[d(x) = \sqrt{(x+2)^{2} +[3+\sqrt{5-(x-1)^{2}}]^{2}}\] Differentiate and set equal to zero \[d'(x) = \frac{3+3\frac{1-x}{\sqrt{5-(x-1)^{2}}}}{d(x)} = 0\] \[\rightarrow x = 1 \pm \sqrt{\frac{5}{2}}\] \[\rightarrow y = 2 \pm \sqrt{\frac{5}{2}}\] nearest point (-.58, .419)...........farthest point (2.58,3.58)